Jelly sells 62 shares of stock she owns for a total of $433. If the stock was in two different companies, one selling at $6.50 a share & the other at $7.25 a share, how many of each did she sell?
Answer by stanbon(75887)

Jelly sells 62 shares of stock she owns for a total of $433. If the stock was in two different companies, one selling at $6.50 a share & the other at $7.25 a share, how many of each did she sell?
----
Quantity Equation:: x + s = 62 shares
Value Equation::: 6.50x+7.25s = 433 dollars
---------------------------
Modify for elimination:
6.5x + 6.5s = 6.5*62
6.5x + 7.25s= 433
-----------------------------
Subtract and solve for "s":
0.75s = 30
-----
s = 40 shares (# of 7.25 shares)
x = 22 shares (# of 6.50 shares)
------------------
Cheers,
Stan H.